OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y = -x + 2 x + 5 = y

hero (hero):

Next time write them on two separate lines: y = -x + 2 y = x + 5

hero (hero):

In this case you can just add the fractions since x's cancel leaving 2y = 7 Divide by 2 to get y = 7/2 Now plug y back into one of the original equations to get x Alternatively, just graph it.

OpenStudy (anonymous):

But how many solutions are there?

hero (hero):

With linear, when the slopes are different, 100% of the time there is only one solution.

OpenStudy (anonymous):

Ok thanks!